# A substellar flyby that shaped the orbits of the giant planets

### v1

@@ 3.1 Best Match
Figure 2: The time evolution of the giant planets before, during, and after the flyby shown in Figure 1. The encounter time is indicated by the vertical line at $t\approx 0.2\,\mathrm{Myr}$. The upper panel shows the inclination of the planets (with respect the the final invariant plane) while the lower panel shows the semi-major axes, the perihelia, aphelia distances. The maximum eccentricity range of each planet is also indicated.
For each flyby simulation, we calculate the power spectrum of the complex eccentricities and inclinations over 20 Myr and compute its value of $\mathcal{M}$ relative to all the solar system power spectra. Of the $5\cdot 10^{4}$ simulated flybys, we found $422$ cases (approximately $1\%$) that lie within the range of the solar system’s own $\mathcal{M}$ values. As we described above, we consider these to be a close match to the orbital architecture of our solar system.
Figure 1 visualizes the encounter that produces the best match to the solar system’s secular architecture according to our metric. In this case, the flyby object has a mass of $m_{\star}=8.27\,\mathrm{M}_{\mathrm{J}}$, a perihelion of $q_{\star}=1.69\,\mathrm{AU}$, a hyperbolic excess velocity of $v_{\infty}=2.69\,\mathrm{km\,s}^{-1}$, and an inclination of $i_{\star}=131^{\circ}$ with respect to the initial common plane of the planets. Figure 2 shows the time evolution of the semi-major axes $a$ of the giant planets along with the perihelia $q$, aphelia $Q$, and inclinations $i$ of the giant planets before, during, and after the flyby encounter. The maximum eccentricities for each planet are indicated on the right.
Figure 3: Power spectra of the complex eccentricities (left column) and complex inclinations (right column). The black line shows the best matching flyby case (as in Figures 1 and 2), while the color-shaded plots show the ensemble of spectra of the solar system taken in 20 Myr–long segments from an ensemble of 5 Gyr–long integrations. The median power at each frequency of the solar system ensemble is shown as a dark coloured line, with lighter shaded regions representing the first three standard deviation percentiles. The median power spectrum of the solar system is very similar to the simulated best case flyby.
Figure 3 shows the power spectra from the FFTs of the complex eccentricities and complex inclinations for the best matching flyby system (black line) overlaid with the ensemble of power spectra of the solar system. In order to see how a particular flyby system compares to the spectral variation present in the true solar system, we determine the spread in the power for each frequency of the FFTs using the reference set of solar system integrations marking the median and the first three standard deviation percentiles. The peaks in these spectra correspond to the frequencies of the secular modes (Laskar, 1988, 1990, 1993; Šidlichovský and Nesvorný, 1996; Laskar, 2003). One can see that the amplitudes of the secular modes of the giant planets have been excited with a single flyby event to values similar to those of the real solar system.
@@
$s_{5}$-- - --- - --- - --- -
$s_{6}$ $-26.35$ $-28.12$ $7.85$ $3.50$
$s_{7}$ $-2.99$ $-3.31$ $8.87$ $8.02$
$s_{8}$ $-0.69$ $-0.61$ $5.81$ $1.20$
Table 1: A comparison of the secular frequencies and amplitudes of the giant planets (SS) and our best matching flyby system that experienced a $m_{\star}=8.27\,\mathrm{M}_{\mathrm{J}}$ encounter at $q_{\star}=1.69\,\mathrm{AU}$, with $v_{\infty}=2.69\,\mathrm{km\,s}^{-1}$. By convention, $s_{5}$ is the null frequency when inclinations are designated with respect to the invariant plane.
Table 1 provides a more quantitative comparison between the present-day Laplace–Lagrange secular modes of the giant planets and the best matching flyby simulation. For the four giant planets, there are four mode frequencies in the complex eccentricities, traditionally denoted as $g_{5},g_{6},g_{7},g_{8}$, and three in the complex inclinations, $s_{6},s_{7},s_{8}$ (Brouwer and Clemence, 1961). Table 1 shows that all the linear secular mode frequencies of the flyby system are a close match with those of the solar system. This follows from the circumstance that the best matching flyby only mildly disturbs the semi-major axes of the planets: by $+2.6\%$, $-0.03\%$, $-3.2\%$, and $+2.6\%$ for Jupiter, Saturn, Uranus, and Neptune respectively. (Recall that the linear mode frequencies depend only upon the semi-major axes and masses.) We also see in Table 1 that the mode amplitudes of the best matching flyby system have magnitudes of a few hundredths in the eccentricity modes and a few thousandths in the inclination modes. While similar in order-of-magnitude to those of the solar system, they differ in detail. The amplitudes of the inclination modes are a closer match than those of the eccentricity modes.
In particular, it has been recognized for some time that it is most challenging to excite the amplitude of the $g_{5}$ mode to values similar to the observed value by any of a number of plausible mechanisms considered in previous studies (Nesvornỳ and Morbidelli, 2012; Griveaud et al., 2024). This is because the $g_{5}$ mode dominates Jupiter’s orbit. As Jupiter has the largest planetary mass in the solar system, it is consequently the most resistant to orbital perturbations. In that light, note that in the best matching flyby case, we find that the amplitude of the $g_{5}$ mode is about $65\%$ that of the present-day solar system. Overall, we find that in $40\%$ of the $422$ closely matching flyby systems, the amplitude of the $g_{5}$ mode is within a factor of two of the solar system’s. This is a notable success of the substellar flybys as a mechanism to explain the secular mode amplitudes.
@@ 3.2 Inner Solar System
Because many of the closely matching systems experience flybys at perihelion distances in the region of the terrestrial planets, we simulate an additional $10^{4}$ circular, co-planar systems. We use the exact same initial conditions for the giant planets and the exact same flyby event as shown in Figs. 1 and 2, but include the terrestrial planets at their current average semi-major axes, also on circular, co-planar orbits but with random phases, as well as the effect of general relativity (Brown and Rein, 2023). The terrestrial planets are initialized on circular coplanar orbits with their present day semi-major axes.
Remarkably, in the majority of cases we roughly reproduce the secular modes and amplitudes of the entire solar system. In these simulations, we find that within the first 20 Myr following the flyby encounter, only $2\%$ of cases result in an ejected planet, with $4\%$ having at least one planet that reaches an eccentricity greater than $0.6$. However, it becomes more difficult to make a specific claim of similarity because there are even more degrees of freedom and larger range of variation, particularly in the power spectra of the terrestrial planets.
3.3 Probabilities
@@
As discussed above, our simulations show that, starting with initially co-planar and circular orbits, approximately $1\%$ of such close flybys result in secular mode amplitudes similar to those of the real solar system. Thus, the overall probability of the solar system acquiring its current dynamical state via a close encounter with a substellar object is roughly $1$ in $10^{4}$ (becoming as likely as $1$ in $10^{3}$).
Figure 4 shows the encounter parameters of all the flyby objects drawn from our cluster parameters, highlighting in blue the close matches to the solar system. See also Figure 6 which shows all flyby parameters. From here, we can see that the closest matching simulations result from flybys are those with substellar masses ($m_{\star}<50\,\mathrm{M}_{\mathrm{J}}$) and pass between the orbits of Mercury and Uranus ($\sim 0.4<q_{\star}<20\,\mathrm{AU}$ from the Sun). For typical velocities within an open star cluster, extremely close flybys ($q_{\star}<50\,\mathrm{AU}$) with stellar masses ($m_{\star}>80\,\mathrm{M}_{\mathrm{J}}$) almost always result in ejected planets or over-excited secular modes. About $60\%$ of our $5\cdot 10^{4}$ simulations had at least one ejected planet in the first $20$ Myr. Stellar mass flybys at greater distances are much less likely to eject a planet, but are also unable to sufficiently excite the eccentricity modes, particularly Jupiter’s eccentricity (Nesvornỳ and Morbidelli, 2012; Griveaud et al., 2024; Brown and Rein, 2022). It may be possible for a stellar mass encounter to enable the same degree of similarity between an initially circular, co-planar system and the solar system. However, such a stellar flyby would need to be moving significantly faster than what is typically found within an open star cluster system, while also passing extremely close, making such cases exceedingly low probability.
@@ 4 Conclusions
In this paper, we explore a new scenario for the origin of the secular mode amplitudes of the giant planets. We show that a single substellar object making a hyperbolic close encounter with the ancient solar system can account for the observed amplitudes in both the eccentricity and inclination secular modes. We find the most favorable encounter to be a substellar object with encounter parameters $m_{\star}=8.27\,\mathrm{M}_{\mathrm{J}}$, $q_{\star}=1.69\,\mathrm{AU}$, $v_{\infty}=2.69\,\mathrm{km\,s}^{-1}$, and $i_{\star}=131^{\circ}$. Our numerical simulations indicate that there is approximately a $1$-in-$10^{3}$ to $1$-in-$10^{4}$ chance of the necessary encounter parameters being realized with random encounters within an open star cluster having properties expected for where the solar system formed. Given that the estimated population of Sun-like stars in the Galaxy is on the order of $10^{10}$ and that stars are commonly formed in open star clusters, the $1$-in-$10^{4}$ chance is not negligible. In other words, we don’t need to look for a needle in a haystack to find a suitable encounter. Additional numerical simulations also demonstrate that the terrestrial planets have a high probability of not only surviving such an encounter but also themselves acquiring secular mode amplitudes similar to those observed today.
@@
Note that there is a very large parameter space of initial conditions to explore in this scenario (velocity, direction, and mass of the flyby object, as well as the orbital phases of all planets) and the outcomes are sensitive to these initial conditions. That we found a substantial fraction of close matches to the solar system shows that this scenario is not only possible but surprisingly efficient. This result holds regardless of the specific measure of closeness adopted (see Appendix A.2 for more details). However, let us acknowledge that identifying with very high precision the one specific set of flyby encounter parameters that reproduces every observed dynamical feature of the solar system is not possible—it would be as unique as the solar system itself.
Nevertheless, further exploration of this scenario is warranted. For example, simulations could include a wider range of initial semi-major axes. This would allow us to connect the present scenario to other earlier phases in planet formation theories including planet-disk interactions. Future investigations might also look into a more detailed study of the effects on the terrestrial planets, as well as the effect of substellar flybys on the dynamical excitation of minor planets in the asteroid belt and the trans-Neptunian belts, and the capture probability of irregular satellites by the giant planets. Our results suggest that substellar flyby encounters should be taken into account when studying the formation and evolution of the Kuiper belt, trans-Neptunian objects (TNOs), and the Sednoids (Pfalzner et al., 2024). Whereas this study focused on finding close matches for our own solar system, a statistical analysis of all flyby outcomes might also reveal new insights into the formation history of extrasolar planetary systems which show different architectures compared to the solar system (Winn and Fabrycky, 2015; Kane and Wittenmyer, 2024).
@@
The solar system ensemble used in this study can be found at Zenodo.4299102. The data generated for this article will be shared on reasonable request to the corresponding author.
@@
A repository containing a portion of the data underlying this article and code for running the simulations and generating the figures will be found at github.com/zyrxvo/. This research was made possible by the open-source projects Jupyter (Kluyver et al., 2016), iPython (Pérez and Granger, 2007), matplotlib (Hunter, 2007; Droettboom et al., 2016), joblib (Joblib Development Team, 2024) and gnu-parallel (Tange, 2023). The author’s open-source projects REBOUND (Rein and Liu, 2012) and AIRBALL (Brown et al., 2024) that enabled this research are also available.

### v2

@@ 3 Results and Discussion
3.1 Close Matches
For each flyby simulation, we calculate the power spectrum of the complex eccentricities and inclinations over 20 Myr and compute its value of $\mathcal{M}$ relative to all the solar system power spectra. Of the $5\cdot 10^{4}$ simulated flybys, we found $422$ cases ($0.844\%$) that lie within the range of the solar system’s own $\mathcal{M}$ values. As we described above, we consider these to be a close match to the orbital architecture of our solar system.
Given that we initialize the pre-flyby circular orbits of the planets at their currently observed osculating semi-major axes, it is unsurprising that the fundamental secular modes of the solar system are prominent in the power spectra of the subset of flyby systems that incur only small changes in the planets’ semi-major axes. In such cases, the fundamental secular mode frequencies remain near the solar system values. However, the amplitudes of the fundamental modes as well as the power in the rest of the Fourier spectrum can vary considerably, depending on the flyby parameters. With the metric $\mathcal{M}$ we are demanding closeness of the entire power spectrum (of the complex eccentricity and complex inclination), not only closeness of the power in the fundamental secular modes. (In Appendix A.4, we describe a flyby case whose fundamental secular modes are closely matched but the full power spectrum is qualitatively and quantitatively very different from that of the solar system.) This closeness of the entire power spectrum is realized in nearly $1\%$ of the flyby cases. In the $\sim 99\%$ of cases that are not close matches we find that the eccentricity modes are significantly under- or over-excited or at least one planet gets ejected in the first 20 Myr.
Figure 2 shows the encounter parameters of all the flyby objects drawn from our stellar cluster parameters, highlighting in blue the close matches to the solar system. (See also Figure A.5 in the Appendix which shows all flyby parameters and their joint distributions in a corner plot.) From here, we can see that the closest matching flybys are those with substellar masses ($m_{\star}<80\,\mathrm{M}_{\mathrm{J}}$), pass between the orbits of Mercury and Uranus ($\sim 0.4\,\mathrm{AU}<q_{\star}<20\,\mathrm{AU}$ from the Sun), and have a hyperbolic excess velocity $v_{\infty}$ of 1–3 $\mathrm{km\,s}^{-1}$. For typical velocities within an open star cluster, extremely close flybys ($q_{\star}<50\,\mathrm{AU}$) with stellar masses ($m_{\star}>80\,\mathrm{M}_{\mathrm{J}}$) almost always result in ejected planets or over-excited secular modes. About $60\%$ of our $5\cdot 10^{4}$ simulations had at least one ejected planet in the first $20$ Myr. Stellar mass flybys at greater distances are much less likely to eject a planet, but are also unable to sufficiently excite the eccentricity modes, particularly Jupiter’s eccentricity (Nesvornỳ and Morbidelli, 2012; Griveaud et al., 2024; Brown and Rein, 2022). It may be possible for a stellar mass encounter to enable the same degree of similarity between an initially circular, coplanar system and the solar system. However, such a stellar flyby would need to be moving significantly faster than what is typically found within an open star cluster system, while also passing extremely close, making such cases exceedingly low probability.
The statistics of the flyby simulation outcomes are summarized as follows. The most probable outcome is a disruption of the planetary system ($\sim 60\%$ of cases). In the next most probable outcome ($\sim 39\%$ of cases) the eccentricities are with roughly equal probability either over or under excited, while the inclinations are consistently under excited when compared to the solar system. In the remaining $0.844\%$ of cases, the outcome is a system similar to the solar system (as per our metric $\mathcal{M}$, Eq. 4).
@@ 3.2 Best Match
Figure 3 visualizes the encounter that produces the best match to the solar system’s secular architecture according to our metric. In this case, the flyby object has a mass of $m_{\star}=8.27\,\mathrm{M}_{\mathrm{J}}$, a perihelion of $q_{\star}=1.69\,\mathrm{AU}$, a hyperbolic excess velocity of $v_{\infty}=2.69\,\mathrm{km\,s}^{-1}$, and an inclination of $i_{\star}=131^{\circ}$ with respect to the initial common plane of the planets. Figure 4 shows the time evolution of the semi-major axes $a$ of the giant planets along with the perihelia $q$, aphelia $Q$, and inclinations $i$ of the giant planets before, during, and after the flyby encounter. The maximum eccentricities for each planet are indicated on the right side of the figure.
@@
Figure 4: The time evolution of the giant planets before, during, and after the flyby shown in Figure 3. The encounter time is indicated by the vertical line at $t\approx 0.2\,\mathrm{Myr}$. The upper panel shows the inclination of the planets (with respect to the final invariant plane) while the lower panel shows the semi-major axes, the perihelia, aphelia distances. The maximum eccentricity range of each planet is also indicated.
[…]
Figure 5 shows the power spectra from the FFTs of the complex eccentricities and complex inclinations for the best matching flyby system (black line) overlaid with the ensemble of power spectra of the solar system. In order to see how a particular flyby system compares to the spectral variation present in the true solar system, we determine the spread in the power for each frequency of the FFTs using the reference set of solar system integrations marking the median and the first three standard deviations. The most prominent peaks in these spectra correspond to the frequencies of the fundamental secular modes (Laskar, 1988, 1990, 1993; Šidlichovský and Nesvorný, 1996; Laskar, 2003). One can see that the amplitudes of the fundamental secular modes of the giant planets have been excited with a single flyby event to values similar to those of the real solar system.
@@
$s_{5}$ -- -- -- --
$s_{6}$ 26.35-26.35 28.12-28.12 $7.85$ $3.50$
$s_{7}$ 2.99-2.99 3.31-3.31 $8.87$ $8.02$
$s_{8}$ 0.69-0.69 0.61-0.61 $5.81$ $1.20$
Table 1: A comparison of the secular frequencies and amplitudes of the giant planets ($S$) and our best matching flyby system that experienced a $m_{\star}=8.27\,\mathrm{M}_{\mathrm{J}}$ encounter at $q_{\star}=1.69\,\mathrm{AU}$, with $v_{\infty}=2.69\,\mathrm{km\,s}^{-1}$. By convention, $s_{5}$ is the null frequency when inclinations are designated with respect to the invariant plane.
Table 1 provides a quantitative comparison between the present-day fundamental secular modes of the giant planets and the best matching flyby simulation. For the four giant planets, there are four fundamental mode frequencies in the complex eccentricities, traditionally denoted as $g_{5},g_{6},g_{7},g_{8}$, and three in the complex inclinations, $s_{6},s_{7},s_{8}$ (Brouwer and Clemence, 1961). Table 1 shows that all the fundamental secular mode frequencies of the flyby system are a close match with those of the solar system. As noted previously, this follows from the circumstance that the best matching flyby only mildly disturbs the semi-major axes of the planets: by $2.6\%$, $0.03\%$, $3.2\%$, and $2.6\%$ for Jupiter, Saturn, Uranus, and Neptune respectively. (Recall that the fundamental mode frequencies depend only upon the semi-major axes and masses.) We also see in Table 1 that the mode amplitudes of the best matching flyby system have magnitudes of a few hundredths in the eccentricity modes and a few thousandths in the inclination modes. While similar in magnitude to those of the solar system, they differ in detail. The amplitudes of the inclination modes are a closer match than those of the eccentricity modes.
In particular, it has been recognized for some time that it is most challenging to excite the amplitude of the $g_{5}$ mode to values similar to the observed value by any of a number of plausible mechanisms considered in previous studies (Nesvornỳ and Morbidelli, 2012; Griveaud et al., 2024). This is because the $g_{5}$ mode dominates Jupiter’s orbit. As Jupiter has the largest planetary mass in the solar system, it is consequently the most resistant to orbital perturbations. In that light, note that in the best matching flyby case, we find that the amplitude of the $g_{5}$ mode is about $65\%$ that of the present-day solar system. Overall, we find that in $40\%$ of the $422$ closely matching flyby systems, the amplitude of the $g_{5}$ mode is within a factor of two of the solar system’s. This is a notable success of substellar flybys as a mechanism to explain the secular mode amplitudes.
@@ 3.3 Inner Solar System
Because many of the closely matching systems experience flybys at perihelion distances in the region of the terrestrial planets, we simulate an additional $10^{4}$ circular, coplanar systems. We use the exact same initial conditions for the giant planets and the exact same flyby event as shown in Figs. 3 and 4, but include the terrestrial planets at their current semi-major axes, also on circular, coplanar orbits but with random phases, as well as the effect of general relativity (Brown and Rein, 2023).
Remarkably, in the majority of cases we roughly reproduce the secular modes and amplitudes of the entire solar system. In these simulations, we find that within the first 20 Myr following the flyby encounter, only $2\%$ of cases result in an ejected planet, with $4\%$ having at least one planet that reaches an eccentricity greater than $0.6$. However, it becomes more difficult to make a specific quantitative claim of similarity for the following reasons.
The inclusion of the terrestrial planets means that there are many more degrees of freedom, and we find that the power spectra in these systems are noisier than the power spectra of the giant planets alone. Consequently, the distribution of the metric values $\mathcal{M}$ has a much larger range. For the solar system compared to itself, the distribution of $\mathcal{M}$ is broadened, shifted, and has a more extended tail relative to its distribution when the inner planets are not included. (The reader is referred to the discussion at the end of Section 2.4 for more context.) Due to the broader distributions of $\mathcal{M}$, many more of our flyby simulations qualify as close matches to the solar system.
3.4 Flyby Probabilities
@@
$\Gamma\approx 1.3\cdot 10^{-4}\,\mathrm{Myr}^{-1}.$ (5)
This encounter rate implies a 1-in-76 chance of a close flyby within $100$ Myr. With the other IMFs discussed in Section 2, we find similar probabilities of a close flyby over $100$ Myr: in the range from 1-in-130 (Chabrier, 2005) to 1-in-71 (Maschberger, 2013).
This probability for close encounters may appear surprisingly high compared with previous works that have investigated stellar flybys of the solar system (Jiménez-Torres et al., 2013; Pfalzner et al., 2018; Pfalzner and Vincke, 2020; Batygin et al., 2020; Brown and Rein, 2022). The reason for this higher probability is the large population of substellar and super-Jovian mass objects included in the IMFs we have considered, whereas previous studies were limited to the stellar IMFs [^1] Li et al. (2025) did look at planetary mass encounters in the context of Hilda asteroids..
As discussed above, our simulations show that, starting with initially coplanar and circular orbits, $0.844\%$ of such close flybys result in power spectra of the planets’ eccentricities and inclinations similar to those of the real solar system. Thus, the overall chance of the solar system planets acquiring their orbital eccentricities and inclinations via a close encounter with a substellar object is 1-in-9005. For the other IMFs we considered, the chance is between 1-in-8412 and 1-in-15403. The probabilities can also be scaled to other stellar environments. For example, if one chooses the lower end of the cluster dispersal timescale of 10 Myr, then the chance decreases by a factor of 10 because we make the simplifying assumtion that the stellar density remains static over the lifetime of the cluster.
@@ 4 Conclusions
In this paper, we explore a new scenario for the origin of the secular mode amplitudes of the giant planets. We show that a single substellar object making a hyperbolic close encounter with the ancient solar system can account for the observed amplitudes in both the eccentricity and inclination fundamental secular modes. We find the most favorable encounter to be a substellar object with encounter parameters $m_{\star}=8.27\,\mathrm{M}_{\mathrm{J}}$, $q_{\star}=1.69\,\mathrm{AU}$, $v_{\infty}=2.69\,\mathrm{km\,s}^{-1}$, and $i_{\star}=131^{\circ}$. Our numerical simulations indicate that there is a 1-in-9005 chance that the necessary encounter parameters can be realized with random encounters within an open star cluster having properties expected for where the solar system formed. Given that the estimated population of Sun-like stars in the Galaxy is on the order of $10^{10}$ and that stars are commonly formed in open star clusters, the $10^{-4}$ probability is not negligible. Even if we choose the most pessimistic scenario, with an IMF that has a much lower fraction of substellar objects (Chabrier, 2005), and a 10 times faster open cluster dispersal timescale, the probability is still $7\cdot 10^{-6}$. In other words, we don’t need to look for a needle in a haystack to find a suitable encounter. Additional numerical simulations also demonstrate that the terrestrial planets have a high probability of not only surviving such an encounter but also themselves acquiring secular mode amplitudes similar to those observed today.
@@
Note that there is a very large parameter space of initial conditions to explore in this scenario (velocity, direction, and mass of the flyby object, as well as the orbital phases of all planets) and the outcomes are sensitive to these initial conditions. That we found a substantial fraction of close matches to the solar system shows that this scenario is not only possible but surprisingly efficient. This result holds regardless of the specific measure of closeness adopted (see Appendix A.2A.4 for more details). However, let us acknowledge that identifying with very high precision the one specific set of flyby encounter parameters that reproduces every observed dynamical feature of the solar system is not possible—it would be as unique as the solar system itself.
Nevertheless, further exploration of this scenario is warranted. For example, simulations could include a wider range of initial semi-major axes. This would allow us to connect the present scenario to other earlier phases in planet formation theories including planet-disk interactions. Future investigations might also look into a more detailed study of the effects on the terrestrial planets, as well as the effect of substellar flybys on the dynamical excitation of minor planets in the asteroid belt and the trans-Neptunian belts, and the capture probability of irregular satellites by the giant planets. Our results suggest that substellar flyby encounters should be taken into account when studying the formation and evolution of the Kuiper belt, trans-Neptunian objects (TNOs), and the Sednoids (Pfalzner et al., 2024).
Whereas this study focused on finding close matches for our own solar system, a statistical analysis of all flyby outcomes might also reveal new insights into the formation history of extrasolar planetary systems which show different architectures compared to the solar system (Winn and Fabrycky, 2015; Kane and Wittenmyer, 2024). To illustrate, we comment on some broader implications of the system disruption statistics in our flyby simulations. We found that $\sim 60\%$ of flyby cases are disruptive, resulting in ejected planets. Combining this result with the estimated flyby encounter rate (Eq. 5), the implication is that a multi-planet system has a 1-in-1280 to 1-in-128 chance of a disruptive flyby over the 10–100 Myr cluster lifetime. This estimate indicates that most planetary systems would avoid disruptive flyby encounters, roughly consistent with previous studies (e.g. Malmberg et al., 2011). We can also estimate that, for the $\mathcal{O}(10^{10})$ Sun-like stars in the Galaxy, the rate of planet ejection by flybys implies $\mathcal{O}(10^{7}-10^{8})$ ejected planets contributing to the Galaxy’s free-floating planet population. These estimates must be taken with caution as they assume that the flyby outcomes have similar statistics for multi-planetary systems with varied planetary masses, initial orbital radii and open cluster properties.
@@
The solar system ensemble used in this study can be found at Zenodo.4299102 and the data generated for this article can be found at Zenodo.15288677.
@@
A repository containing a portion of the data underlying this article and code for running the simulations and generating the figures can be found at github.com/zyrxvo/Secular-Origin. This research was made possible by the open-source projects Jupyter (Kluyver et al., 2016), iPython (Pérez and Granger, 2007), matplotlib (Hunter, 2007; Droettboom et al., 2016), joblib (Joblib Development Team, 2024) and gnu-parallel (Tange, 2023). The open-source projects REBOUND (Rein and Liu, 2012) and AIRBALL (Brown et al., 2024) that enabled this research are also available.